OpenStudy (anonymous):

One number is 8 times another number. The numbers are both positive and have a difference of 70. Find the numbers. Which of the following equations could be used to solve the problem? x - 8x = 70 8 x - x = 70 8 - x = 70 x - 8 = 70

OpenStudy (anonymous):

10 and 80.

OpenStudy (anonymous):

Because 10 times 8 is 80. And 80 - 10 is 70.

OpenStudy (anonymous):

x - 8x = 70 8 x - x = 70 8 - x = 70 x - 8 = 70

OpenStudy (anonymous):

And the second equation could be used!

OpenStudy (anonymous):

Look: 8x-x=70 8x - x is 7x So 7x= 70 Divide by 7 on both sides. x = 10 So therefore one of the numbers is ten, and once you know that, you just do plus 70, so 80, and you have your answers!